In 2000, the population of Israel was approximately 6.04 million and by 2050 it is projected to grown to 10 million. Use the exponential growth model A = Aoe^kt, in which t is the number of years after 2000, to find an exponential growth function that models the data.
In which year will Israel's population be 9 million?

Answer by stanbon(48535)

In 2000, the population of Israel was approximately 6.04 million and by 2050 it is projected to grown to 10 million. Use the exponential growth model A = Aoe^kt, in which t is the number of years after 2000, to find an exponential growth function that models the data.
----------------------
You have two points: (0,6.04) and (50,10)
A(t) = Aoe^kt
If t = 50
10 = 6.04e^(10k)
1.656 = e^(10k)
Take the ln of both sides to get:
ln(1.656) = 10k
k = 0.05042
EQUATION is Therefore A(t) = 6.04e^(0.05042t)
----------------
In which year will Israel's population be 9 million?
9 = 6.04e^(0.05042t)
1.49 = e^(0.05042t)
Take the ln of both sides to get:
0.05042t = 0.3988206
t = 7.91 years
---------
Nine million in the year 2008
===============
Cheers,
Stan H.